Suppose you had an 8%, $10,000 semi-annual bond with three years remaining to maturity.The yield on new three-year bonds of comparable quality is 6%. Calculate what your bond is worth in the secondary market.
Answer #1

Information provided:

Par value= future value= $10,000

Coupon rate= 8%/2= 4%

Coupon payment= 0.04*10,000= $400

Interest rate= 6%/2= 3% per semi-annual period

Time= 3 years*2= 6 semi-annual periods

The present value is calculated to compute the worth of the bond.

Enter the below in a financial calculator to compute the present value:

FV= 10,000

I/Y= 3

N= 6

PMT= 400

Press the CPT and PV to compute the present value.

The value obtained is 10,541.72.

Therefore, the worth of the bond is $10,541.72.

In case of any query, kindly comment on the solution.